VIBRATIONAL SPECTRA AND STRUCTURES OF METAL (Ti, Ni, Pd) MONONITROSYL COMPLEXES. AN IR MATRIX ISOLATION AND DFT STUDY

\underline{LAHOUARI KRIM}, M. ESMAIL ALIKHANI and LAURENT MANCERON, LAPDIR/Spectroschimie Moléculaire, UMR7075 CNRS-P6, Université Pierre et Marie Curie, Bat F74, Boîte 49, 4 place Jussieu, 75252 Paris, Cedex 05, France.